Find the x and y intercept for 5y-x=10 I need help
Answer by nerdybill(7384)

Find the x and y intercept for 5y-x=10
.
Rearrange the given equation into the "point-slope" form:
y = mx + b
where
m is the slope
b is the y-intercept at (0,b)
.
5y-x=10
5y=x+10
y = (1/5)x + 2
From inspection, then we have
slope is 1/5
y-intercept is at (0,2)